OpenStudy (anonymous):

The width of a table is 8 feet less than 4 times its length. Identify the tables length and width if its area is 32 ft to the second power.

OpenStudy (anonymous):

\[ w=4l-8\\ wl=32\\ (4l-8)\cdot l=32\\ 4l^2-8l-32=0\\ l^2-2l-8=0\\ (l-4)(l+2)=0 \]Can you finish it from there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok i got 4 and 8

OpenStudy (anonymous):

To find out if that's right, plug it back in and see if it works. If the length is 4, and the width is 8, then it is true that the width is 8 less than 4 times the length, and it is also true that the area is 32. Voila. :)

OpenStudy (anonymous):

awesome thanks